[Brahms-l] Shift report 20050512 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Thu May 12 2005 - 16:10:01 EDT
Shifter leader was SB
Shifters were 


1. Beam for about 40 minutes. Machine development since 9:30 AM. No physics running yet. Problem with yellow beam loss does not seem to
have been resolved so far.
2.  FS @2.3 deg @3450A, MRS @90deg @2255A
--------------------------------------------
run14600: trig2(17k),trig3(5.4),trig4(2.8),trig5(26k),trig6(25k)
run14601: trig2(27k),trig3(8.4k),trig4(4.1k),trig5(38k),trig6(40k)
-----------------------------------------------------
FFS @3 deg BFS @2.3 deg @3450A, MRS @90deg @2255A
-----------------------------------------------------
run14602: trig2(13k),trig3(21k),trig4(11k),trig5(63k),trig6(26k)
-------------------------------------------------------------

3. No detector problems

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Thu May 12 16:10:21 2005

This archive was generated by hypermail 2.1.8 : Thu May 12 2005 - 16:10:37 EDT